OpenStudy (anonymous):

simplify e^(-2x)/(3e^(-4x))

OpenStudy (whpalmer4):

\[\frac{e^{-2x}}{3e^{-4x}}\]Correct?

OpenStudy (anonymous):

can it be e^2x/3

OpenStudy (whpalmer4):

it can be!

OpenStudy (whpalmer4):

Or if you don't like fractions, \[e^{2x}3^{-1}\]but that's maybe a bit unusual...

OpenStudy (anonymous):

ok cool i just checked my calculator it rights

OpenStudy (whpalmer4):

Did you do it by flipping the fraction and then simplifying, or did you just subtract the exponents right away?

OpenStudy (anonymous):

both lol

OpenStudy (whpalmer4):

Ah, good job! I like someone who checks their work.

OpenStudy (anonymous):

i was curious of the 3e with the log rules

OpenStudy (anonymous):

then i just realized its just a constant

OpenStudy (whpalmer4):

Yep, a very special constant, but a constant nonetheless. Have you learned about imaginary/complex numbers yet?

OpenStudy (anonymous):

im in diff q

OpenStudy (whpalmer4):

Okay, then hopefully yes :-) I imagine you've seen \[e^{\pi i} - 1 = 0\]

OpenStudy (whpalmer4):

except that should be a plus sign, duh

OpenStudy (anonymous):

this was just to see if the website was legit and i was kind of iffy on the subject....but the website really works

OpenStudy (anonymous):

yea i have seen it

OpenStudy (anonymous):

so the eq e^xi +1=0